OpenStudy (anonymous):

Use the Quadratic Formula to solve the quadratic equation. 9t^2 − 4t + 2 = 0

OpenStudy (saifoo.khan):

\[t = \frac{-(-2) \pm \sqrt{(-4)^2-4(9)(2)}}{2(9)}\]

OpenStudy (anonymous):

it was looking for solutions sorry.

OpenStudy (anonymous):

p.s. how do you create the fractional bar?

OpenStudy (saifoo.khan):

right click on the photo, hit "show source".

OpenStudy (saifoo.khan):

\[\frac 19 (2-i \sqrt{14}),\frac 19 (2+i \sqrt{14}),\]

hero (hero):

Saifoo, the form of your solution, as posted, sucks

OpenStudy (saifoo.khan):

Good to hear. thanks for the views.

hero (hero):

If I were a teacher, I would take off a quarter of a point for posting a solution that way

OpenStudy (saifoo.khan):

But you aren't. ;)

hero (hero):

And neither are you ;)

OpenStudy (saifoo.khan):

I never said I'm.

hero (hero):

You use I'm in the wrong context

OpenStudy (saifoo.khan):

your statement makes no sense.

hero (hero):

You're not qualified to make such a statement

OpenStudy (anonymous):

saifoo i really need help with the problem i have just posted

OpenStudy (saifoo.khan):

Alright.